A student’s test scores are 86, 72, 85, and 90. What is the lowest score that
the student can get on the next test and still have a test average (mean) of at
least 80?

Answer by edjones(7311)

Let x=the next test score to give an average of 80
(86+72+85+90+x)/5=80
(86+72+85+90+x)=400
333+x=400
x=67
.
Ed